OpenStudy (anonymous):

Using l'hopitals rule please compute:

OpenStudy (anonymous):

\[\lim_{n \rightarrow \infty} n^2(1-\cos(\frac{ 5.2 }{ n}))\]

OpenStudy (anonymous):

@tkhunny

OpenStudy (anonymous):

@ikram002p

OpenStudy (tkhunny):

For starters, we need to get it into a form where the rule applies. Which indeterminate form are we looking at?

OpenStudy (anonymous):

It doesn't mention in the question. Can you give me the easiest one?

OpenStudy (tkhunny):

?? It's not a worked example. It's a think about it problem. Can we use l'hospital or not? It has to be an indeterminate form or it's no good.

OpenStudy (anonymous):

Yeah ofcourse we can use l'hopitals if we divide by n which would make it n^-2

OpenStudy (tkhunny):

You didn't quite say that right, but it seems you have the right idea. Good work. So, now we have this: \(\dfrac{1 - \cos\left(\dfrac{5.2}{n}\right)}{\dfrac{1}{n^{2}}}\) What indetermiante form is that?

OpenStudy (anonymous):

0/inf

OpenStudy (anonymous):

sorry thats 0/0

OpenStudy (tkhunny):

Since the first one is no good, I am delighted that you fixed it. Okay, now for the derivatives.

OpenStudy (anonymous):

the top would be sin(5.2/n) (-5.2n^-2) all divided by -2n^-3

OpenStudy (anonymous):

@tkhunny are you there?

OpenStudy (anonymous):

Its here where I got stuck and needed help.

OpenStudy (tkhunny):

Yes. I'm waiting for you to simplify your result. That's horrid.

OpenStudy (anonymous):

I can't do it, I got stuck over here.

OpenStudy (ipwnbunnies):

*eats popcorn* This is great. Hint: Get rid of the denominator!

OpenStudy (tkhunny):

I actually don't believe that. Did you take algebra or not? I'm going with "yes". The numerator has n^(-2) and the denominator has n^(-3). Simplify that pair.

OpenStudy (ipwnbunnies):

Got stuck on the easiest part mate lol.

OpenStudy (anonymous):

thats n^(-1) on bottom which is equivalent to n on the numerator. I have been doing all types of integrals all day! Sorry guys <3.

OpenStudy (tkhunny):

Well, take a deep breath and do one more. Keep doing it until you get \(2.6n\sin(5.2/n)\)

OpenStudy (anonymous):

After that what do I do? The answer isn't infinity, its 13.52. So I take the derivative again?

OpenStudy (tkhunny):

So jumpy. Did you take that deep breath? You MAY need another derivative. Did anyone tell you that l'hospotal's rule doesn't always work, even if it is an indeterminate form? Anyway, start over. Is this expression an indeterminate form? Just like we started - think it over at each step. As soon as it's NOT in indeterminate form, we have to think of something else. So, it it? CAN we do another derivative?

OpenStudy (anonymous):

Yeah, use the product rule. to find the derivative. I know that you can take the limit inside the function but I don't know when to do that, when we have got the right form?

OpenStudy (tkhunny):

?? That doesn't make any sense. Did you look at this form? \(\dfrac{2.6\cdot\sin(5.2/n)}{1/n}\) Is that an indeterminate form?

OpenStudy (anonymous):

Yes this is inf/0

OpenStudy (tkhunny):

That is not correct. Look at the numerator a little harder. Remember, it's sine up there, not cosine.

OpenStudy (anonymous):

0/0 then!

OpenStudy (anonymous):

@iPwnBunnies are you still eating popcorn mate?

OpenStudy (ipwnbunnies):

Yep. Now, apply L'Hopital's Rule again.

OpenStudy (tkhunny):

Please do.

OpenStudy (anonymous):

2.6n = u and sin(5.2/n) = v u' = 2.6 and v' =cos(5.2/n)(1/n) Product rule = u'v + v'u 2.6sin(5.2/n) + cos(5.2/n)(1/n)(2.6n)

OpenStudy (tkhunny):

There is no product rule. The variable appears only once in both numerator and denominator.

OpenStudy (anonymous):

I took the n back up and included it with the 2.6. It is equivalent to what you wrote.

OpenStudy (tkhunny):

l'hospital's rule does not tell you to find the derivative of the entire expression. You cannot do it. You must find the indeterminate form and then treat numerator and denominator separately.

OpenStudy (anonymous):

Oh okay! Then here is the part that I went wrong in. So it would be 2.6cos(5.2/n)(1/n) all divided by -1/n^2

OpenStudy (anonymous):

@tkhunny am I right or did I do a silly mistake again?

OpenStudy (tkhunny):

There are no silly mistakes when they cause you to learn. Did you make any deliberate errors? Try the numerator again.

OpenStudy (anonymous):

@tkhunny No but its late night and I have been studying all day I can't think anymore. Anyways I got your idea, thank you very much for your help! I really appreciate it.

OpenStudy (tkhunny):

If you are careful, you should get \(13.52\cos(5.2/n)\) and that should be helpful.

OpenStudy (ipwnbunnies):

Aww man, you're almost there. You messed up the chain rule in the numerator.

OpenStudy (anonymous):

@tkhunny thank you very much for your help!

OpenStudy (anonymous):

@iPwnBunnies Thank you very much for your sensitive emotions.